Lithium chloride has a solubility of 55 gg of LiClLiCl in 100 gg of H2OH2O at 25 ∘C∘C. Determine if each of the following mixtures forms an unsaturated or saturated solution at 25 ∘C∘C:

answer the following questions with the information given above.

1.) adding 10 gg of LiClLiCl to 15 gg of H2O

2.) adding 23 gg of LiClLiCl to 50. gg of H2O

3.) adding 73 gg of LiClLiCl to 150. gg of H2O
Answer #1

solubility of LICl = 55 g / 100 g = 0.55 g / g water

1.) adding 10 g of LiClto 15 g of H2O:

excess undissolved LiCl = 10 - (0.55 * 15) = 1.75 g

thus

this is saturated solution.

2.) adding 23 g of LiCl to 50. g of H2O

50 g H2O can soluble the 55 * 50 / 100 = 27.5 g LiCl maximum.

but here 23 g LiCl only.

so

this is unsaturated solution.

3.) adding 73 g of LiCl to 150. g of H2O

150. g of H2O dissolve the maximum = 55 * 150 / 100 = 82.5 g LiCl

but there are 73 g LiCl only.

so

this is unsaturated solution.